OpenStudy (anonymous):

how many 4 person committees can be chosen from a group of 12 people?

OpenStudy (anonymous):

Use the equation of nCr=n!/r!(n-r)! so 12C4=

OpenStudy (anonymous):

thanks..but i needed help with the work.

OpenStudy (anonymous):

Okay so 12C4= 12!/4!(12-4)! Then, 12C4 = 12!/4!(8!) so 12C4 = 12*11*10*9/4! Because the 8! cancels out.

OpenStudy (anonymous):

thank you soo much

OpenStudy (anonymous):

Does that all make sense?

OpenStudy (anonymous):

yes it does. :)